### v1

# Ceci N’est Pas un Drone: Investigating the Impact of Design Representation on Design Decision Making When Using GenAI

Abstract
With generative AI-powered design tools, designers and engineers can efficiently generate large numbers of design ideas. However, efficient exploration of these ideas requires designers to select a smaller group of potential solutions for further development. Therefore, the ability to judge and evaluate designs is critical for the successful use of generative design tools. Different design representation modalities can potentially affect designers’ judgments. This work investigates how different design modalities, including visual rendering, numerical performance data, and a combination of both, affect designers’ design selections from AI-generated design concepts for Uncrewed Aerial Vehicles. We found that different design modalities do affect designers’ choices. Unexpectedly, we found that providing only numerical design performance data can lead to the best ability to select optimal designs. We also found that participants prefer visually conventional designs with axis-symmetry. The findings of this work provide insights into the interaction between human users and generative design systems.
1 Introduction
Engineering design is a complex and creative exercise soria_zurita_survey_2017; song_toward_2020; song_when_2021; xu_adaptation_2024; xu_mind_2025 that involves innovative ideation, critical thinking, technical analysis, and iterative problem solving. It also requires designers to make tradeoffs in their search for solutions, balancing between exploration and exploitation in the design space march_exploration_1991; tabeau_exploration_2017; okamoto_proposal_2022. Exploration often requires designers to look for solutions that are original and novel, exemplified by design activities like ideation and brainstorming march_exploration_1991; gupta_interplay_2006. In contrast, exploitation emphasizes the refinement of existing details and making choices towards greater efficiency; this mindset often becomes important during design validation and iterative prototyping march_exploration_1991; gupta_interplay_2006. Artificial Intelligence (AI) technologies present new opportunities for engineers and designers to achieve improved productivity in both exploration and exploitation song_when_2021; xu_adaptation_2024.
Generative AI can substantially support design exploration by rapidly providing an abundance of feasible design solutions, usually significantly more than what would be possible with human effort alone, further highlighting the importance of the subsequent design selection and exploitation process to materialize those innovations oh_deep_2019; zhu_generative_2022; zhu_generative_2023; kim_effect_2023. Engineers and designers, restricted by limited time and resources, will need to use their expertise and judgment to select a smaller group of candidate designs for further validation and prototyping wallace_methods_1995; gembarski_making_2021. Therefore, the ability to judge the quality of designs, especially when presented with numerous options, is critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025.
Prior studies have shown that the modality in which a design is presented affects people’s perceptions reid_impact_2013; detchprohm_effect_2025; barnawal_evaluation_2017; schulze-meesen_impact_2023; derya_ozcelik_buskermolen_effect_2015, which could potentially affect their judgments. It is common for designers to use visualizations of their designs, such as visual renderings (geometric representations of physical products) or sketches, to aid communications and collaborations (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021). Reid et al. reid_impact_2013 found that different visual representations affect customer subjective preferences and objective measurements of the products, but not their judgments on product function attributes. Also, Detchprohm et al. detchprohm_effect_2025 showed in a study that the quality of visual representations of the designs will not affect the perceived functionality of the product. Studies have also shown that visual representations facilitate better communication and feedback compared to textual information barnawal_evaluation_2017; schulze-meesen_impact_2023. However, results on the differences between visual representations are mixed. One study by Barnawal et al. barnawal_evaluation_2017 suggests that 3D design representations do not have a significant advantage over 2D ones for communicating design concepts, but have higher usability, while another study by Buskermolen et al. derya_ozcelik_buskermolen_effect_2015 found that the motion (stills vs. an animation) and visual quality do not affect concept comprehension, but visual quality can affect the style of communication. Padilla et al. padilla_decision_2018 investigated the effect of different data visualizations on people’s judgments and decision-making when interpreting data, and found that data visualization affects decision-making by invoking different cognitive processes. The complex nature of engineering design tasks requires both objective information delivery and subjective evaluation of the feasibility of design concepts outside of design data. The potential biases introduced by different design representation modalities could weaken designers’ judgment abilities and, therefore, negatively impact the usefulness of generative AI systems in engineering design.
It seems an easy conclusion that humans rely on visualization techniques to do design well, given the ubiquity of design visualization in design practices. Studies have shown that design visualization (e.g., sketches) adds additional information to numerical performance data as a cognitive tool ullman_importance_1990; goldschmidt_dialectics_1991; suwa_what_1997, suggesting that designers have heuristics that enable them to extract non-textual information from design visualization. However, does that necessarily mean designers can make better judgments and design selections with design visualization? Making the design selection process more challenging, AI-generated designs can be "weird" and "unusual", with aesthetics that humans may find unattractive despite performance advantages loos_towards_2022. In the context of engineering design, this raises further questions: Do engineers actually need to see the design visualization to make good judgments and design choices? Or is the numerical performance data, along with proper data visualization, all they need? Do design visualization and geometric rendering carry any additional value to the design decision-making process outside of what design numerical performance data can offer? Research in education has shown that visualization deepens learners’ understanding of the subject and improves learning outcomes as an epistemic object evagorou_role_2015; schoenherr_learning_2024. It is possible that the presence of design visualization and design renderings enables human heuristics and allows engineers and designers to capture features and information that are missed by objective functions or mere numerical performance data. It may be critical to have design renderings and visualization presented for engineers to make accurate judgments. Nevertheless, it is also possible that the presence of design visualization introduces unwanted bias or design fixations, leading to sub-optimal design decisions.
That leads to our central research question: Do different design modalities affect human decision-making behavior and their ability to make optimal selections when presented with AI-generated design solutions? So far, there is insufficient evidence in the design and AI research community to answer this question or to explicitly and systematically examine the impact of design modalities on design decision-making. We hypothesize that different design modalities can affect engineers’ decision-making. We also hypothesize that using a design visualization (geometric representations of physical products) as a representation modality can negatively impact engineers’ ability to identify optimal and novel design solutions.
To test our hypotheses, we designed and conducted a within-subjects experiment across three studies targeting different populations to study whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to choose the optimal designs in a list of AI-generated design ideas. These three studies progressively build on one another to address emergent phenomena observed. During a single study, the participants individually considered three UAV (uncrewed aerial vehicles) design problems, where they were asked to select the most optimal drone design from the provided list of design options and briefly explain their choice. For the first study, we targeted drone hobbyists and drone pilots at a U.S. research university. During the first study, we observed that the number of design options provided could affect the participants’ choices. To further investigate the issue, we conducted the second study with STEM students at a U.S. research university through an engineering design related course. For the third study, we recruited aerospace engineers and designers from a large aerospace company in order to examine potential expert-novice differences. Combined, these experiments investigate whether and how different types of design representations affect engineers’ decision-making when working with generative-AI design tools.
Our results confirm both of our hypotheses. We found that different design modalities do affect engineers’ decision-making when using AI-powered generative design tools. Specifically, in this study, designers provided with numerical design performance data only lead to the best ability to select optimal designs, while only seeing design renderings provides marginal help in selecting optimal designs. Unexpectedly, when participants saw both the numerical design performance data and the design rendering, they performed more poorly compared to when they saw the numerical design performance data alone. However, we found that the participants prefer the best-performing designs as long as those designs possess traditional and symmetrical appearances. We also found in this study that the number of design options provided affects people’s ability to choose optimal designs, and a large number of design options can overwhelm engineers and lead to suboptimal design choices.
This paper contributes empirical evidence on how design representation modalities affect designers’ decisions when using AI-powered generative design tools, specifically through:
  • Empirical findings from three studies with drone hobbyists, STEM students, and aerospace professionals, showing that different representation modalities (visual rendering, numerical data, and visual rendering + numerical data) affect design choices with generative design tools.
  • Evidence suggesting that presenting only numerical performance data leads to the most accurate identification of optimal designs, while adding visual renderings can reduce accuracy. Also, a larger number of design choices overwhelms participants, reducing selection accuracy.
  • Analysis of human heuristics and preferences, revealing that designers prefer conventional, symmetrical designs with good performance.
2 Related Work
2.1 AI in engineering design
In an effort to facilitate solving complex engineering design problems, the engineering design research community has been studying automated design tools since the 1980s maher_hi-rise_1985; smithers_-based_1989. With the rapid development of modern AI and Machine Learning (ML) technologies, a wide variety of studies have investigated the implementation of AI and ML in the engineering design process, including design exploration and concept generation kim_ai_2019; raina_learning_2019; camburn_computer-aided_2020; valdez_framework_2021; zhu_generative_2022; zhu_generative_2023; kim_effect_2023; saadi_generative_2023; khanolkar_mapping_2023; joosten_comparing_2024, design concept evaluation camburn_machine_2020; song_decoding_2022; demirel_human-centered_2024, design optimization sharpe_comparative_2019; nie_topologygan_2021; behzadi_gantl_2021; senhora_machine_2022; wang_generative_2023; maze_diffusion_2023, and prototyping and manufacturing dering_unsupervised_2017; williams_design_2019; qin_research_2022; tercan_machine_2022; kumar_machine_2023. Those implementations of AI in engineering design are believed to improve the efficiency of the design process mirhoseini_graph_2021; yuksel_review_2023, and can even improve the quality of the design solutions joosten_comparing_2024.
More noticeably, a recent focus of the research community is the adoption of AI-powered design tools for design generation oh_deep_2019; chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. AI-powered generative design tools can promptly create a large number of design solutions chen_inverse_2021; heyrani_nobari_creativegan_2021; regenwetter_deep_2022. With the help of AI-powered generative design tools, engineers and designers can search for potential design solutions more efficiently with a larger scope and at a lower cost koch_design_2017; oh_deep_2019; camburn_computer-aided_2020; camburn_machine_2020; dering_unsupervised_2017; byrne_application_2025. In general, generative AI is believed to enhance the design ideation and concept generation process oh_deep_2019; kim_effect_2023. However, the abundance of potential design solutions may pose new challenges for designers when selecting optimal designs.
2.2 Design decision making and judgment ability: key for successful AI implementation
Concept selection and design decision making are an essential part of engineering design wallace_methods_1995; gembarski_making_2021. Restricted by limited time and resources, engineers and designers constantly face trade-offs and design decision-making in the engineering design process division_on_engineering_and_physical_sciences_theoretical_2001; otto_trade-off_1991; nickel_contextual_2024. In order to make informed decisions, designers need to consider different factors and features, and seek a balance between them kalsi_comprehensive_1999; division_on_engineering_and_physical_sciences_theoretical_2001. As a result, the judgment ability and the ability to evaluate designs are critical for the successful implementation of generative design tools hong_generative_2023; chen_how_2025; fang_generative_2025. Experienced engineers and designers can better leverage their domain knowledge to prioritize promising AI suggestions and better utilize this new technology tambe_reskilling_2025. It is important to consider the impact on the judgment abilities of the designers and domain experts when introducing design tools.
2.3 Design representation modalities and the impact on judgment ability
Design representations are frequently used in engineering design ideation, communication, and collaboration (xu_mind_2025; henderson_flexible_1991). Studies have shown that design representations, especially visual design representations, can serve as boundary objects for effective information exchange within design teams, facilitating collaboration and mitigating misunderstanding (xu_mind_2025; bucciarelli_between_2002; subrahmanian_boundary_2003; kalay_enhancing_2001).
Among the common design modalities, numerical performance data is an intuitive and straightforward approach to conveying design information in an accurate and concise manner, especially when paired with appropriate data visualization abi_akle_graphical_2015; araci_trade-off_2017; cibulski_paved_2020. However, constrained by its textual nature and narrow representation of design features, design information delivered by numerical performance data could be limited in scope (e.g., incomplete or fragmented information on design development and justification cheng_role_2019; mirabito_feature_2024).
In comparison, design visualization (geometric representations of physical products) is another common design modality that is visual-based and more intuitive, and has a more inclusive representation of physical design features (larkin_why_1987; henderson_flexible_1991; ivanov_visualization_2024). Visual design representations contain rich information about the design, especially for the structure of the design, allowing for accurate and comprehensive interpretation (xu_mind_2025; tsai_how_2017). As a result, design visualizations are widely used in different stages of the engineering design process (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021; veisz_computer-aided_2012). It is common practice in the engineering design industry to use visualization of the designs, such as visual renderings or sketches, to aid communications and collaborations (henderson_flexible_1991; haggman_connections_2015; tsai_how_2017; atit_shah_learners_2021). Research has shown the positive effects of using visualization for engineering design practices, including facilitating ideation, communication, and collaboration, and improving shared understanding (suwa_what_1996; mckoy_influence_2001; tversky_what_2002; tversky_sketches_2003; heiser_sketches_2004; macomber_role_2011; worinkeng_can_2013; xu_mind_2025).
However, research has also cautioned the potential risks of overreliance on design visualizations, including design fixation and detriments to design creativity jansson_design_1991; atilola_representing_2015; atilola_effects_2016; amann_fixation_1988; viswanathan_design_2013. These effects can affect designers’ judgment, and hinder their willingness and ability to explore novel solutions suggested by generative AI systems. Furthermore, studies have shown that design modalities could affect people’s perceptions of the design, and, therefore, affect their judgments reid_impact_2013; derya_ozcelik_buskermolen_effect_2015; barnawal_evaluation_2017; schulze-meesen_impact_2023; detchprohm_effect_2025. Judgment biases can also result from the use of particular modalities padilla_decision_2018. All these potential hindrances can weaken designers’ abilities to objectively evaluate designs and select optimal design solutions, and, therefore, negatively impact the development and implementation of generative engineering design tools.
3 Methodology
To test our hypotheses, we designed and conducted a within-subjects experiment across 3 populations to study whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to choose the optimal designs in a list of AI-generated design ideas.
3.1 Participants
In total, 160 participants were recruited, with 156 college students and 4 aerospace industry professionals. Out of the 156 recruited college students, 29 of them participated in study 1 as self-identified drone pilots and drone hobbyists from a U.S. research university. The other 127 of the college student participants in study 2 are from a sophomore-level engineering design course in the Department of Mechanical Engineering at the same U.S. research university. Participation was voluntary, and the participants were compensated with a \$10 Amazon gift card. Participants recruited from the design course are also compensated with course credits. The 4 aerospace industry participants in study 3 were recruited from a major aerospace company. The ethnicity, age, and gender of the participants did not affect the recruitment process. All participants were over the age of 18 when recruited.
  • Among the 29 drone pilots and drone hobbyists participants in study 1, 11 identified as women, and 18 identified as men. Also, 7 drone hobbyist participants identified as White, 1 identified as Black or African American, 17 identified as Asian, 3 identified as other, and 1 preferred not to disclose their ethnicity.
  • Among the 127 STEM student participants in study 2, 62 identified as women, 57 identified as men, 2 identified as non-binary or third gender, 1 identified as other, and 5 preferred not to identify themselves. Also, 55 STEM student participants identified as White, 8 identified as Black or African American, 35 identified as Asian, 1 identified as Native Hawaiian or Pacific Islander, 20 identified as other, and 8 preferred not to disclose their ethnicity.
  • Among the 4 professional participants in study 3, all 4 of them identified as men. Also, 3 professional participants identified as White, and 1 identified as Native Hawaiian or Pacific Islander.
3.2 Study Design
The university ethics review board approves human-subjects research, and they approved this project. Each of the three studies consisted of one online survey. In the survey, participants individually consider three UAV (uncrewed aerial vehicles) design problems. The design problems are similar in terms of difficulty and design scope, but with different design requirements (e.g., low cost versus low maintenance). A detailed problem description and design requirements for each design problem are provided in Appendix A.
The UAV designs used in the survey are from the AircraftVerse dataset developed by SRI International and Southwest Research Institute cobb_aircraftverse_2023. The UAV designs in this dataset are created using a proprietary generative AI system developed by SRI International cobb_aircraftverse_2023 with real-world, off-the-shelf components with accurate physical properties. All the UAV designs used in this study are simulated for flight performance using a simulator developed for this work with the help of SRI International and Southwest Research Institute. We then documented the performance of each UAV design when conducting the tasks described in the design problems, enabling us to numerically determine the level of optimality of the design solutions presented to the participants.
The participants are asked to consider three design problems about UAVs with different design requirements. The three design problems are accompanied by three sets of design solutions. For each problem, they are tasked with selecting the best drone design and are also allowed to share the reasoning for their selection. The number of design solution options is 2, 8, and 16 for the three design problems, respectively. The 4-option case is used in training. The three sets of design solutions are the same for each experimental condition, but with different modalities, and are presented in random order. For each design problem, the three sets of design solutions are each presented by either 1) visual design rendering, 2) numerical design performance data with data visualization, or 3) visual design rendering plus numerical performance data and data visualization. The visual design rendering is presented with interactive 3D renderings of the UAV designs. The numerical design performance data is presented as tables, along with spider plots, which is a common practice in the aerospace industry. Figure 1 shows an example drone design using these two design modalities, as presented to the participants. The performance data available to the participants are max hover time, max travel distance, mass, max air speed, battery voltage, total cost, and max lift. Not all of the listed performance metrics are relevant to every design problem.
(a) Example design modality - visual rendering.
(b) Example design modality - numerical data with data visualization.
Figure 1: Examples of design modalities used in the studies.
The main identifying features and key performance metrics of the drone designs used in this study are shown in Tables LABEL:tab:drone_feature_1, LABEL:tab:drone_feature_2, and LABEL:tab:drone_feature_3 in Appendix B. We characterize each drone according to several identifying features; drone designs can be axis-symmetric or non-axisymmetric, planar or non-planar, and one-plane or off-plane. Some drones are axis-symmetric with respect to the center of the main body, with propellers aligned axisymmetrically to one another around the center of the main body of the drones on one horizontal plane. Non-axisymmetric designs are ones in which the propellers are not axisymmetric to each other around the center of the main body of the drones. Non-planar means that not all the propellers are on a single horizontal plane. Off-plane means that at least one propeller is on a different horizontal plane than the main body of the drone. We define designs that are axis-symmetric and one-plane as conventional designs, and designs that have atypical features (i.e., non-axisymmetric, non-planar, or off-plane) as unusual designs. Additionally, we introduce a measure Rod Length to Propeller Area Ratio which is the ratio between the total length of connecting rods of the drone and the total effective propeller area, with a unit of $m^{-1}$.
Key performance metrics include Maximum Thrust, Effective Lift, Hover Time, and Total Cost. Maximum Thrust is the maximum instant thrust that the drone can provide, based on our simulation, in Newtons ($N$). It is one of the most important features in evaluating the performance of the drone designs under the design requirements of the design problems, as it determines not only the amount of weight the drone can lift but also how quickly the drone can move. Effective Lift is the maximum weight ($kg$) that the drone can carry, determined by the maximum lift minus the drone’s own weight. Hover Time describes the maximum time that the drone can stay in the air and maintain the target position with minimal deviation. Total Cost is the total cost of the drone.
Neither Rod Length to Propeller Area Ratio nor Maximum Thrust is provided to the participants during the study. The former was introduced during analysis to support interpretation, while the latter was replaced with Max Lift to provide a more readily-interpretable metric of payload capacity.
The participant groups and design question orders of all three studies are shown in Table 1 below. Each study is detailed in the following subsections.
3.2.1 Study 1
For the first study, we targeted drone hobbyists and drone pilots at a U.S. research university. In total, 29 student drone hobbyists were recruited. In this study, the three design problems are presented in the order of 2-option, 8-option, and 16-option.
3.2.2 Study 2
In the first study (Study 1), we observed that the number of options could affect the participants’ choices, and that we see a drastically different trend with different design problems, especially those with 8 design options versus those with 2 or 16 options. We conducted the second study to identify whether the irregularity is due to the design problem itself or to the order in which the design problems were shown. We recruited 127 STEM students at the U.S. research university through an engineering design related course. The second study has two different survey design conditions. The first condition, Study 2A, presents the design problems in the order of 2-option, 8-option, and 16-option, identical to the first study. The second version of the survey, Study 2B, presents the design problems in the order of 2-option, 16-option, and 8-option. Both versions of the survey are identical in all aspects other than the order in which the design problems are shown. The design problems and the design options are the same as those used in the first study.
3.2.3 Study 3
Since the first two studies focused on the college student, we seek to examine potential differences between that novice population and experienced engineering professionals working in the aerospace industry. Thus, we ran a third study with participants who were working as aerospace engineers and designers in a large aerospace company. Since we did not find the order in which the design problems are presented to affect participants’ abilities to pick the better designs ($F(1,1135)=0.351,p=0.553$) in the second study, we kept the survey design identical to the first study, with the three design problems with 2, 8, and 16 options, in order.
Study Number Participant Group
(Number of Participants)
Design Question Order
(Number of Design Options for Each Design Problem)
Study 1 Drone Hobbyists (29) 2 Options -> 8 Options -> 16 Options
Study 2A STEM students (63) 2 Options -> 8 Options -> 16 Options
Study 2B STEM students (64) 2 Options -> 16 Options -> 8 Options
Study 3 Professionals (4) 2 Options -> 8 Options -> 16 Options
Table 1: Study Designs
3.3 Procedure
The participants in all three studies completed the study using an online survey tool (Qualtrics). After giving consent, the participants begin the study and are instructed to consider three design problems about UAVs with different design requirements. For each design problem, the design options are shown in three different types of design representation modalities, namely, visual rendering, numerical performance data with data visualization, and visual rendering with numerical performance data and data visualization. The participants are informed that the designs are automatically generated by an AI system and that we need their expertise to evaluate and determine the feasibility of these solutions. The participants are also informed that the designs are tested in an advanced simulator, but they still need to utilize their engineering experience and knowledge to evaluate the designs, considering real-world scenarios. The study then asks the participant to choose the most optimal drone design from the provided list of design options (see Appendix B) with respect to several design restrictions. For each design selection, the survey also asks the participants to briefly explain their choice in the form of a textual response.
3.4 Measurement and Data Analysis
In this section, we outline the metrics and analytical procedures used to evaluate participants’ decision-making across experimental conditions. The analysis incorporates different quantitative measures, capturing the variability and optimality of participants’ design selections as well as the thematic content of their written explanations. Together, these measures provide a comprehensive view of participants’ performance, enabling assessment of both the outcomes of their choices and the underlying reasoning processes.
3.4.1 Multiple Choice Question Response Entropy
We first calculated the Shannon entropy of the choices for each design problem. A smaller entropy value corresponds to greater order and less randomness in the participants’ choices, suggesting the participants are making similar decisions instead of random choices. The entropy is calculated as
$H=-\sum_{i=1}^{n}p_{i}\ln p_{i}$ (1)
where H is the Shannon entropy and $p_{i}$ is the probability of the $i$-th design being chosen. We estimate these probabilities simply as
$p_{i}=\frac{\text{\# times design }i\text{ was chosen}}{\text{total choices made}}$ (2)
3.4.2 Multiple Choice Question Response Accuracy
We also want to examine how different design modalities affect the participants’ ability to select better designs. In order to consistently select the best design, we employ the Pareto-TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution) method hwang_multiple_1981; hwang_new_1993; hu_new_2023. Pareto-TOPSIS ranks candidate options based on their weighted distance to the ideal and the worst options, and is widely used in engineering design and HCI research as an objective means of establishing optimality hu_new_2023; alizadeh_data-driven_2019; souaille_interactive_2022; bertoni_iterative_2019; wang_enhanced_2021; chatterjee_probabilistic_2017. The option closest to the ideal and farthest from the worst is deemed the most optimal. In this work, we applied the same weight to all the design objectives in the design problems, namely, Hover Time, Maximum Thrust, Effective Lift, and Total Cost.
The Pareto-TOPSIS optimal solutions for problems with 2, 8, and 16 options are design 18393, design 20155, and design 20985, respectively. The Pareto-TOPSIS optimal designs (dominating Pareto optimal and best TOPSIS) are indicated with an asterisk in Tables LABEL:tab:drone_rendering_1-LABEL:tab:drone_feature_3 in Appendix B. We then calculated the accuracy of participants selecting the Pareto-TOPSIS optimal designs and performed ANOVA (Analysis of Variance) tests on the accuracies to unveil differences across experimental conditions.
3.4.3 Multiple Choice Question Response Change Across Modalities
We seek to examine how the participants picked different designs under different design representation modalities, and specifically how those selections changed between problems. We illustrated and analyzed the changes in participants’ design choices across design representation modalities with Sankey diagrams for each experiment.
3.4.4 Textual Response
Participants’ textual justification for their design choices is analyzed to better understand their decision making. We coded each participant’s textual responses based on the main driving feature(s) cited for selecting their chosen design. A code was only assigned if the participant explicitly mentioned the term or a near synonym (including different word forms).
We conclude 4 main driving features (non-mutually exclusive) for the visual rendering only condition, including symmetry, stability, reliability, and cost. These four factors are common contributing "visual inference" factors, where the participant is making their judgment based on the visualization of the design, including information inferred from the visualization. Please note that here cost is also inferred from the visual renderings as no numerical performance data is provided in this condition.
As for the numerical performance data only condition, we identified five main driving features (non-mutually exclusive), including hover time, lift, cost, stability, and reliability. Here, information about hover time, lift, and cost is provided with numerical data, and the participants also managed to infer information about stability, reliability of the design from the numerical data as well.
The main driving features (non-mutually exclusive) for the visual rendering with numerical performance data condition include symmetry, stability, reliability, hover time, lift, and cost. Here, the participants can infer design information from both the visual renderings and the provided performance metrics. Therefore, we categorize symmetry, stability, and reliability as visual inferences, as these factors are more likely to be inferred from the visual rendering. And, we categorize hover time, lift, and cost as performance metrics, as these factors are more likely to be obtained from the numerical performance data.
We then calculated and analyzed the frequency of those driving factors that appear in textual responses to find the main driving factors for participants’ decision-making.
4 Results
The results of each study will be shown separately in the following sub-sections, followed by the results of the analysis on the text responses. Each study subsection includes quantitative analysis on the multiple-choice question responses, including response entropy, accuracy, and design choice change across modalities.
4.1 Study 1 - Drone Hobbyists
The participants in Study 1 are drone hobbyists. The number of participants that have chosen each design is shown in Figures 2, 3, and 4, with detailed distribution available in Appendix C, Table 12, 13, and 14.
Figure 2: Study 1: Participants’ choice transition in Problem 1 - 2 options.
Figure 3: Study 1: Participants’ choice transition in Problem 2 - 8 options.
Figure 4: Study 1: Participants’ choice transition in Problem 3 - 16 options.
4.1.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 5 below. Overall, the entropy values are relatively small, indicating low randomness in the participants’ answers, suggesting the participants are making informed decisions based on the provided design information.
Figure 5: Study 1: Entropy of participants’ design choices
4.1.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 6. Most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Design modalities did affect participants’ decision-making here. However, there are no consistent changes found when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The transitions of participants’ design choice changes are shown in Figures 2, 3, and 4.
Figure 6: Study 1: Participants’ choice of design for consecutive questions in three design problems. V is with visual rendering, D is with numerical performance data, and M is with both visual rendering and numerical performance data.
Participants’ selection of conventional design (axis-symmetric and one-plane) vs unusual design (non-axisymmetric, non-planar, or off-plane) is shown in Figure 7. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, when transitioning from numerical performance data to the mixed modality (both visual rendering and numerical performance data).
Figure 7: Study 1: Participants’ choice for conventional design vs unusual design in percentage.
4.1.3 Multiple Choice Question Response Accuracy
The average results of drone hobbyist participants’ accuracy on selecting the optimal designs are shown in Figure 8, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is listed in Appendix C, Table 15.
Figure 8: Study 1: Participants’ accuracy on selecting the optimal designs. Error bars represent a 95% confidence interval. Participants are drone hobbyists.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem 1 with 2 options: $F(2,84)=31.207$, $p<0.001$, $\eta_{p}^{2}=0.426$. Design problem 2 with 8 options: $F(2,84)=8.809$, $p<0.001$, $\eta_{p}^{2}=0.173$. Design problem 3 with 16 options: $F(2,84)=6.774$, $p=0.002$, $\eta_{p}^{2}=0.139$). The results indicate that the design modality does affect participants’ abilities to choose the better designs.
In addition, we performed two-way ANOVA tests on the design modality and the number of design options (design problem) to see if the number of design options provided would change participants’ decision-making results (Effect of Design Problem: $F(2,252)=26.605,p<0.001$; Interaction Effect: $F(4,252)=3.268,p=0.012$). Both the design modality and the number of design options affect the accuracy of participants, and the interaction effect between the design modality and the design problem is also significant.
Since different design problems result in different design choices, we looked into the differences caused by the design modality separately for each design problem. Specifically, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 2. In more than half of the pairwise comparisons, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between the numerical data and mixed modality conditions, and between the visual rendering and mixed modality conditions. However, there are no significant differences found between numerical data and the mixed modality conditions for picking the optimal designs in the design problem with 16 design options. There are also no significant differences found between visual rendering and the mixed modality conditions for picking the optimal designs in both the design problem with 8 design options and the design problem with 16 design options.
Comparison
Group
Design
Problem
T-test results
Visual vs Data 2 options t(28) = 10.360, p < 0.001, Cohen’s d = 1.924
8 options t(28) = 3.923, p < 0.001, Cohen’s d = 0.728
16 options t(28) = 3.839, p < 0.001, Cohen’s d = 0.713
Data vs Visual + Data 2 options t(28) = 3.839, p < 0.001, Cohen’s d = 0.713
8 options t(28) = 2.816, p =0.009, Cohen’s d = 0.523
16 options t(28) = 1.983, p = 0.057, Cohen’s d = 0.368
Visual vs Visual + Data 2 options t(28) = 4.218, p < 0.001, Cohen’s d = 0.783
8 options t(28) = 1.684, p = 0.103, Cohen’s d = 0.313
16 options t(28) = 2.415, p = 0.023, Cohen’s d = 0.448
Table 2: Study 1: T-test results on the effect of design modalities on the accuracy of participants in selecting the optimal designs.
4.2 Study 2 - Engineering Students
Study 1 demonstrated that the number of options affects participants’ choices, and we observed a different trend in the problem with 8 design options compared to the problems with 2 or 16 options. To determine whether the irregularity is significantly influenced by the order in which the design problems were presented, we conducted the second study. The participants in Study 2A and Study 2B are engineering students at the aforementioned U.S. research university. The design problems for Study 2A are presented in the order of 2-option, 8-option, and 16-option, and those in Study 2B are presented in the order of 2-option, 16-option, and 8-option.
The results from Study 2A and Study 2B are first compared based on the accuracy of selecting the optimal designs. We performed two-way ANOVA tests on the design modality and the order in which the design questions are presented to see if the order of design questions would affect participants’ choices. The results of the two-way ANOVA tests are shown in Table 3 below. The order in which the design questions were presented did not significantly affect the accuracy of participants.
Design
Problem
Effect of Question Order Interaction Effect
2 options F(1, 375) = 0.172, p = 0.678 F(2, 375) = 0.322, p = 0.725
8 options F(1, 375) = 0.011, p = 0.917 F(2, 375) = 0.035, p = 0.965
16 options F(1, 375) = 2.235, p = 0.136 F(2, 375) = 1.924, p = 0.147
Table 3: Study 2: Two-way ANOVA test results on the interaction effect between the design modalities and the order in which the design questions are presented, for Study 2A and Study 2B. Results are based on the accuracy of participants in selecting the optimal designs.
Further, we performed an Extra‑Sum‑of‑Squares F–test on the participants’ accuracy to determine if the order of design questions affects participants’ overall responses. We compared two nested regression models, with the reduced model using "modality" and "question list" as predictors, and the full model using an additional predictor, "order of question list." The results suggest that there is no evidence that the additional variable (i.e., "order of question list") adds predictive power ($F(1,1135)=0.351,p=0.553$). The patterns of responses collected from Study 2A and Study 2B are not statistically different. Therefore, the results of Study 2A and Study 2B will be combined as one study and presented together. The number of participants that have chosen each design is shown in Figures 9, 10, and 11, with detailed distribution available in Appendix C, Tables 16, 17, and 18.
Figure 9: Study 2: Participants’ choice transition in Problem 1 - 2 options.
Figure 10: Study 2: Participants’ choice transition in Problem 2 - 8 options.
Figure 11: Study 2: Participants’ choice transition in Problem 3 - 16 options.
4.2.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem and for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 12 below. Overall, the entropy values are relatively small, indicating low randomness in the participants’ responses, suggesting the participants are making informed decisions based on the provided design information.
Figure 12: Study 2: Entropy of participants’ design choices
4.2.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 13. Most participants’ choices changed when the design modality changed from visual rendering to numerical performance data. Thus, design modalities did affect participants’ decision-making here. However, changes are inconsistent when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The changes in participants’ design choices are shown in Figures 9, 10, and 11.
Figure 13: Study 2: Participants’ choice of design for consecutive questions in three design problems. V is with visual rendering, D is with numerical performance data, and M is with both visual rendering and numerical performance data.
Participants’ selection of conventional designs (axis-symmetric and one-plane) versus unusual designs (non-axis-symmetric, non-planar, or off-plane) is shown in Figure 14. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked unusual designs when visual renderings became available (i.e., when transitioning from numerical performance data to the mixed modality).
Figure 14: Study 2: Participants’ choice for conventional design vs unusual design in percentage.
4.2.3 Multiple Choice Question Response Accuracy
The average results of engineering student participants’ accuracy on selecting the optimal designs are shown in Figure 15, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is available in Appendix C, Table 19.
Figure 15: Study 2: Participants’ accuracy on selecting the optimal designs. Error bars represent 95% confidence interval. Participants are STEM students.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem 1 with 2 options: $F(2,378)=218.552$, $p<0.001$, $\eta_{p}^{2}=0.536$. Design problem 2 with 8 options: $F(2,378)=44.940$, $p<0.001$, $\eta_{p}^{2}=0.192$. Design problem 3 with 16 options: $F(2,378)=38.645$, $p<0.001$, $\eta_{p}^{2}=0.170$). The results indicate that the design modality would affect participants’ abilities to choose the "better" designs.
Additionally, we performed two-way ANOVA tests on the design modality and the number of design options (design problem) to see if the number of design options provided would change participants’ decision-making results (Effect of Design Problem: $F(2,1134)=95.540,p<0.001$. Interaction Effect: $F(4,1134)=14.233,p<0.001$). Both the design modality and the number of design options affect the accuracy of participants, and the interaction effect between the design modality and the design problem is also significant.
Since different design problems result in different participants’ choices, we looked into the differences caused by the design modality separately for each design problem. Specifically, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 4. In general, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between numerical data and mixed modality conditions, and between visual rendering and mixed modality conditions.
Comparison
Group
Design
Problem
T-test results
Visual vs Data 2 options t(126) = 26.071, p < 0.001, Cohen’s d = 2.313
8 options t(126) = 10.140, p < 0.001, Cohen’s d = 0.900
16 options t(126) = 9.358, p < 0.001, Cohen’s d = 0.830
Data vs Visual + Data 2 options t(126) = 7.110, p < 0.001, Cohen’s d = 0.631
8 options t(126) = 3.405, p < 0.001, Cohen’s d = 0.302
16 options t(126) = 3.415, p < 0.001, Cohen’s d = 0.303
Visual vs Visual + Data 2 options t(126) = 11.902, p < 0.001, Cohen’s d = 1.056
8 options t(126) = 7.624, p < 0.001, Cohen’s d = 0.677
16 options t(126) = 6.394, p < 0.001, Cohen’s d = 0.567
Table 4: Study 2: T-test results on the effect of design modalities on the accuracy of participants in selecting the optimal designs.
4.3 Study 3 - Aerospace Professionals
The participants in Study 3 are engineers and design professionals working in a large aerospace company. No statistical tests were performed on the data collected from Study 3, as the number of participants is notably smaller than what is required to satisfy the normality assumption of statistical tests. The number of participants that have chosen each design is shown in Figures 16, 17, and 18, with detailed distribution available in Appendix C, Tables  20, 21, and 22.
Figure 16: Study 3: Participants’ choice transition in Problem 1 - 2 options.
Figure 17: Study 3: Participants’ choice transition in Problem 2 - 8 options.
Figure 18: Study 3: Participants’ choice transition in Problem 3 - 16 options.
4.3.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 19 below. Overall, the entropy values are small, indicating low randomness in the participants’ answers, suggesting the participants are making informed decisions based on the provided design information.
Figure 19: Study 3: Entropy of participants’ design choices
4.3.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 20. Most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Interestingly, most participants’ design choices also changed when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. Design modalities did affect participants’ decision-making here. The transitions of participants’ design choice changes are shown in Figures  16, 17, and 18.
Figure 20: Study 3: Participants’ choice of design for consecutive questions in three design problems. V is with visual rendering, D is with numerical performance data, and M is with both visual rendering and numerical performance data.
Participants’ design selection of conventional design (axis-symmetric and one-plane) vs unusual design (non-axisymmetric, non-planar, or off-plane) is shown in Figure 21. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, when transitioning from numerical performance data to the mixed modality (both visual rendering and numerical performance data).
Figure 21: Study 3: Participants’ choice for conventional design vs unusual design in percentage.
4.3.3 Multiple Choice Question Response Accuracy
The average results of aerospace professional participants’ accuracy on selecting the optimal designs are shown in Figure 22, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is shown in Appendix C, Table 23.
Figure 22: Study 3: Participants’ accuracy on selecting the optimal designs. Error bars represent 95% confidence interval. Participants are aerospace professionals.
4.4 Textual Response
The frequencies of the driving factors for participants’ decision-making that appear in textual responses are shown in Figure 23, 24, and 25 below, for the three different design representation modality conditions.
Figure 23: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the visual rendering condition. The data is shown in the format of "number of responses (percentage)".
Figure 24: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the numerical performance data condition. The data is shown in the format of "number of responses (percentage)".
Figure 25: The frequencies of the driving features for participants’ decision-making that appear in textual responses for the visual rendering + numerical performance data condition. The data is shown in the format of "number of responses (percentage)".
Overall, the results here demonstrated that the participants considered different factors when presented with different design representation modalities, and they frequently considered these driving factors when evaluating the designs. From the results of the frequencies of the driving factors for participants’ decision-making that appear in textual responses, we gained insight into the designers’ heuristics for their design choices.
When only visual renderings are provided, designers can effectively utilize their design heuristics and infer the real-world performance of the drone designs, mainly based on symmetry, stability, reliability, and cost of the design. Such a judgment is logical as symmetry, stability, and reliability can hint at the flight capabilities of the drone when no numerical performance data is provided. Interestingly, many participants in this study also estimated and compared the cost of the designs based on the visual rendering, a factor we thought only feasible with numerical performance data.
In comparison, when only numerical performance data is provided, participants focused on hover time, lift, and cost of the designs. These are also the factors that are most important for the flight task described in the design problem. Surprisingly, some participants also managed to infer information about stability, reliability of the design from the numerical data as well.
We also found that more participants considered the performance metrics than the visual inferences when both visual renderings and numerical performance data are available. The commonly considered design factors and features are similar to the previous two conditions, which include symmetry, stability, reliability, hover time, lift, and cost. This might suggest the participants in this study prioritize performance metrics over visual inferences when both are provided.
5 Discussion
5.1 Design modalities affect design decision making
We found that different design modalities appear to affect engineers’ decision-making, and that only using a visual design rendering as a design modality can negatively impact engineers’ ability to identify the optimal design solutions, when using AI-powered generative design tools.
The results of the entropy tests suggest that all three tested modalities can provide some extent of design information, and participants are making informed design decisions, resulting in non-random selections and non-uniform distributions. From the distribution of participants’ answers, there are differences between showing visual design rendering, numerical performance data, and visual design rendering with numerical performance data. The design modality affects participants’ design decision-making process and leads to different results.
The difference between different design modalities is more noticeable when looking at the participants’ choice of design for consecutive questions. For all three participant groups, most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Despite the inconsistency across design problems, a significant portion of participants’ design choices also changed when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. Since the design options are the same across all three modalities, the changes in participants’ choice when facing different design modalities indicate that the design modalities did affect participants’ decision-making. The participants change their minds and possibly select based on different factors when presented with different design modalities, even though the options are the same.
The difference between design modalities is also noticeable when looking at the participants’ ability to select the optimal designs. The three design modalities showed significantly different levels of benefits for selecting the Pareto-TOPSIS optimal design.
Overall, the numerical performance data of the designs gives the most helpful and accurate information for finding the optimal design options. The participants were able to find the optimal designs with much higher accuracy when shown only numerical performance data, compared to the other two modalities. This also shows people’s ability to read and interpret design information from data and spider plots, even though the spider plots face some criticism, including inconsistent areas and shapes caused by axis ordering, misleading area size, and deceptive importance of irrelevant options due to dimension normalization, from the data visualization research community feldman_filled_2013; heijungs_two_2022; duan_origami_2023; abeynayake_efficacy_2023.
The results also suggest that visual renderings of the designs alone might provide the least helpful design information for determining the optimality of the given designs. Participants’ accuracy in selecting optimal designs is quite low. It might suggest that the participants gained little information from the visual rendering of the design alone, or that the visual renderings alone provide distorted design information and lead the participants to inaccurate conclusions (e.g., the propellers can appear smaller if the body of the UAV is larger). However, an alternative explanation is that the participants gained additional information from the visual renderings that are not captured by the performance data and our simulator (e.g., a design might not be easily serviceable, or it might not perform well in cross-winds). Therefore, the difference in accuracy in selecting the optimal designs for the visual rendering condition might suggest that engineers and designers can utilize human heuristics and their design experiences to gain insights that are not easily captured by current computational tools.
Interestingly, participant accuracy decreased when both the visual rendering and the numerical performance data were provided, compared to the numerical performance data only condition. It seems that the addition of visual information changed the participants’ minds, and they then made selections based on other factors. Participants’ textual responses suggest that they are considering and utilizing factors not explicitly shown with the numerical data. For example, in design problem 2 with 8 options, one of the aerospace professional participants (Study 3, P#01) in this study recognized the Pareto-TOPSIS optimal design based on the numerical performance data, but also pointed out the potential design flaws and infeasibility for real-life operations of that design. They then chose the designs with high reliability, stability, and practicality, which were factors they inferred from the visual rendering. Similar to the findings in the visual rendering condition, it is possible that the participants are getting additional information from the rendering that isn’t captured by the performance data. It could also mean the visual renderings are biasing people’s perceptions or causing design fixations. Future work should investigate further whether the addition of visual rendering is positively or negatively impacting designers’ decision-making.
5.2 Designers prefer conventional designs with good performance
In this study, participants showed clear preferences for conventional designs. Drone designs with conventional layouts and features have a significantly higher pick rate, namely, designs that are axis-symmetric and in which all propellers and the main body sit on one plane. Across all three participant pools, the participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, after transitioning from numerical performance data to the mixed modality.
Moreover, there is a significant difference in pick rates for designs with conventional features (axisymmetric and one-plane) versus designs with atypical features (non-axisymmetric, non-planar, or off-plane) for the visual rendering only condition ($t(24)=3.103,p=0.014,\text{Cohen's }d=1.458$). Participants picked more conventional axisymmetric and one-plane designs significantly more often than the other designs. Furthermore, the axisymmetricity seems to be the dominant factor, as there is a significant difference in pick rates for axisymmetric designs versus non-axisymmetric designs for the visual rendering only condition ($t(24)=2.512,p=0.023,\text{Cohen's }d=0.866$). Participants prefer axisymmetric designs. In addition, we found that a Rod Length to Propeller Area Ratio of 5 is a critical divider for participants’ preferences. The pick rates for designs with Rod Length to Propeller Area Ratio greater than 5 are significantly higher than those with Rod Length to Propeller Area Ratio less than 5 ($t(24)=2.840,p=0.017,\text{Cohen's }d=1.209$).
Looking at the changes in participants’ choice of design when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data, a noticeable number of participants moved away from the "unorthodox" designs with better performance, to pick designs with more conventional design features. An example is in design problem 1, where 55 out of 155 participants who chose the "unorthodox" design (design 18393) when shown only numerical performance data moved away and chose the more conventional design 16875 when presented with both visual rendering and numerical performance data (Figure 26). Similarly, 21 and 14 out of 38 and 66 participants, who chose the more unusual design 16763 and design 20985, two of the most picked designs, when provided only numerical performance data, chose the more conventional design 15317 instead when presented with both visual rendering and numerical performance data (Figure 27). These changes in participants’ choice of design when the design modality changed from numerical performance data to visual rendering + numerical performance data clearly show participants’ preference for conventional design features, despite the "unorthodox" designs having better performance data.
Figure 26: Changes in participants’ choice of design for Design Problem 1 when the design modality changed from numerical performance data to visual rendering + numerical performance data.
Figure 27: Changes in participants’ choice of design for Design Problem 3 when the design modality changed from numerical performance data to visual rendering + numerical performance data.
Among the designs with conventional features, design options with better performance have a higher pick rate when numerical performance data is provided. The participants’ accuracy in selecting the Pareto optimal designs within conventional designs is shown in Figures 28, 29, and 30 below (design problem 1 with two options is ruled out because it has only one conventional design option). Overall, the accuracy is high when numerical performance data is provided, in both numerical data and mixed modality conditions, suggesting that participants can identify the better-performing designs among those with conventional features.
Figure 28: Study 1: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are drone hobbyists.
Figure 29: Study 2: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are STEM students.
Figure 30: Study 3: Participants’ accuracy on selecting the optimal designs, when selecting within conventional designs. Error bars represent 95% confidence interval. Participants are aerospace professionals.
Participants’ textual responses for their reasons for the design choices also reflect their preferences for axis-symmetric and "reasonable-looking" designs. One participant in Study 2A said they “feel that symmetrical designs have a big impact on the hover time, so asymmetrical design D, design E, and design H were immediately ruled out” [Study 2A, P#57]. A participant in Study 2A also ruled out asymmetrical designs as “options have large asymmetry would make maintenance harder and more expensive due to specialized parts”[Study 2A, P#19]. Participant #54 in Study 2B explained their design choice as “the rendering seems reasonable”. Another participant explained their preference for symmetrical design, as “symmetrical design would make manufacturing easier” [Study 2A, P#26]. Reliability can be another factor that drives the preference for symmetry, as one participant said “Design D combines symmetry with rotors close to the center of mass to create a reliable design” [Study 2A, P#18].
One participant in Study 2B clearly showed their preferences for "reasonable-looking" and good-performing design in their textual response, as they state “among the designs that look reasonable, Design K has the best statistics with a lower cost, and would probably work best for this scenario” [Study 2B, P#27]. They also doubted the designs with unconventional features despite the better performance data, saying “while Design C has the best statistics, I don’t think it would actually work in real life” [Study 2B, P#27]. Interestingly, there are also participants who have higher faith in the numerical data, “even though A looks strange, it outperforms B in every metric. There is a chance that its unorthodox design lends it some unique advantages” [Study 2B, P#08].
Overall, participants prioritize designs with conventional features, and they prefer the better-performing designs among those with conventional features. Such design heuristics can be helpful for participants to identify the good performers, considering both simulation results and real-life scenarios. However, it can also mean that the participants are unknowingly or unintentionally omitting good-performing designs with atypical design features, leading to potential design fixation.
However, it is important to keep in mind that the performance data and the simulator used in this work may be inaccurate. Some real-life factors can be neglected, leading to an inaccurate evaluation of the drone design performance in real-life scenarios. The participants, however, may have considered these factors, which may have informed their design choice. This is further discussed in the limitations and future works subsection. Also, it is possible that the participants are aware of this imperfection during the studies, and therefore have less faith in the numerical performance data, leading to the preferences for conventional designs seen here.
5.3 Designer’s judgment ability is weakened when presented with a large number of options
From the two-way ANOVA test results on the interaction effect between the design modalities and the number of design options, we found that the number of design options provided can affect the participants’ accuracy in identifying the optimal designs. Since all three design problems have a single most optimal design (based on Pareto-TOPSIS), we compared the participants’ accuracy in selecting the optimal design across design problems with different numbers of design options. We found that, in all three modality conditions, the participants’ accuracy in selecting the optimal design decreases in the design problems with 8 or 16 design options, compared to the problems with only 2 options. This may suggest that with more potential design solutions provided, participants’ ability to determine the optimal design is weakened.
One possibility is that with fewer options shown, participants’ mental capacity is large enough to examine solutions and compare them in pairs throughout and simultaneously, leading to more comprehensive design interpretations. When a large number of candidate solutions are provided, e.g., with 8 or 16 options, the amount of design information that needs to be considered exceeds the mental capacity of the participants. The participants might be overwhelmed with potential options and the pair-wise comparisons when facing complex and challenging design requirements. Their abilities and willingness to explore more designs or more novel designs can be hindered. This finding resonates with research in working memory, which suggests that the capacity of human working memory is limited at a given time and is typically limited to 4 objects for visual working memory and short-term memory luck_capacity_1997; cowan_magical_2001, although the classic research in working memory by George Armitage Miller argued a larger 7 objects with a range of plus or minus 2 miller_magical_1956. Future research should investigate this issue further, along with a more explicit examination of the impact of the number of design options offered on the designer’s decision-making.
5.4 Aerospace professionals might have better judgment abilities
Despite the small group size, the aerospace professionals who participated in this study might have shown a slight tendency toward superior ability to identify better designs.
Comparing the accuracy of selecting the optimal designs in the numerical performance data condition, the aerospace professionals have a higher accuracy on the design question with 8 design options, compared to the student population. Also, the aerospace professionals in this study showed a slightly stronger ability to identify the better design with only numerical performance data, on the design question with 2 design options. Interestingly, their ability to identify the better design with mixed modality (visual rendering + numerical data) is much inferior in comparison, on all three design questions.
Limited by the small sample size, we cannot conclusively determine the impact of participants’ technical and work backgrounds. However, our results might suggest that the technical and work backgrounds could affect participants’ preferences and decision-making when facing AI-generated designs. Future studies should further investigate the effect of technical background and field experience on designers’ ability to identify better designs.
Figure 31: Participants’ accuracy on selecting the optimal designs across participant groups for the design question with 2 options. Error bars represent 95% confidence interval.
Figure 32: Participants’ accuracy on selecting the optimal designs across participant groups for the design question with 8 options. Error bars represent 95% confidence interval.
Figure 33: Participants’ accuracy on selecting the optimal designs across participant groups for the design question with 16 options. Error bars represent 95% confidence interval.
5.5 Limitations and future works
This work faces several limitations. Firstly, the optimality of the design solutions in this work is evaluated with a physics-based simulator. Such a simulator is accurate regarding the physics phenomena and drone design features that are incorporated in the simulator. However, since no simulator considers all factors and scenarios that may occur in real life, many important factors can be neglected in the simulation results for real-life applications. Such imperfection can lead to an inaccurate evaluation of the drone design performance in real-life scenarios. Therefore, the simulation results used in this work and the performance evaluated for each drone design may not be perfectly accurate under the tasks described in the design problems. That means our evaluation of participants’ ability to select the optimal designs can be distorted. It is possible that certain participants conducted a more thorough and realistic evaluation of the designs than the simulator, but their design choice is deemed inferior in this study. This might lead to a seemingly poorer ability to select the optimal designs, while in reality, it is quite the opposite.
Secondly, this study only examined a very specific design problem with relatively high requirements for domain knowledge. Participants with prior UAV design, manufacturing, or flying experience will likely have a more comprehensive heuristic and can demonstrate a stronger ability to evaluate the UAV designs compared to those who don’t, potentially affecting the results. Future work should further investigate the impact of design modalities on more general design problems and other technical domains.
Thirdly, most participants in this study are current college students. Though it is a representative group to demonstrate the impact of different design modalities and experiment conditions, the participants may lack the field experience for some of the realistic technical challenges reflected in the design problems, which sometimes can only be obtained through years of industry-specific work experience.
Another limitation in this work lies in the relatively short length of the design decision-making process simulated here. Even though no hard limit was imposed, most participants finished the study within 40 minutes. Such a time span may not fully and accurately represent the extensive nature of real-world design challenges.
Furthermore, the numerical design performance data is presented along with spider plots, which are regarded as common practice in the aerospace industry. Since extensive research has been done to show that different styles of data visualization can affect the interpretation of visual information, future work should also examine the impact of different design visualization methods on design decision-making perer_integrating_2008; correll_value-suppressing_2018; moritz_formalizing_2019.
Moreover, in this work, we observed that the number of design options provided in the study can alter participants’ ability to identify the optimal designs. It is worth investigating the impact of the number of design options on people’s design decision-making more directly in future studies. Besides, this work focuses more on the design decision-making results rather than the design decision-making process. The heuristics and reasoning behind the decisions made in the work under different conditions are briefly examined. Future work should further and more comprehensively study the influence of different design modalities on designers’ decision-making behavior and rationale.
6 Conclusion
This work examines the impact of different design modalities on the design decision-making process of engineers and designers. More specifically, we investigate whether design visualization affects engineers’ and designers’ ability to select the optimal designs from a list of AI-generated design ideas. We found that different design modalities do affect engineers’ decision-making when using AI-powered generative design tools. For the participants in this study, providing only the numerical design performance data leads to the best accuracy in selecting the most optimal design, while only seeing design renderings provides marginal help in selecting optimal designs. We found that presenting both the numerical design performance data and the design rendering results in worse accuracy compared to seeing the numerical performance data alone, suggesting that engineers can leverage design heuristics when providing the design visualization, or, alternatively, the presence of design visualization can induce design biases and fixations. In addition, we found that the participants generally prefer the best-performing designs as long as those designs possess traditional and symmetrical appearances. Also, we note that the number of design options provided affects people’s ability to choose optimal designs in this study, and a large number of design options can overwhelm engineers and lead to suboptimal design choices. This work deepens our understanding of how people interact with generative AI design systems, and can guide the future development and implementation of design comparison tools used with generative AI systems.
Acknowledgments
Thanks are due to Dominik Moritz and Adam Perer for their indispensable assistance and feedback on the experimental design of this work. The authors are grateful to researchers at SRI International Research Lab and Southwest Research Institute for their generous resources and help in developing the UAV flight simulator. The authors are also grateful to Allison Fisher and Ranald Engelbeck for their invaluable feedback and tremendous help in recruiting participants to make this work possible.
Appendix A Design Problem
Imagine you are a lead design engineer working in an engineering consulting company designing UAVs tailored to customer needs. Your team uses an AI-powered automatic design generation system to help ideate and create initial design solutions. These solutions have been tested in a newly developed computer simulation environment. Your simulation teams assure you that the simulator is the best in the business and that the simulated performance data is accurate. However, as an engineer, you still need to use your expertise and engineering knowledge to inspect and evaluate the designs while considering real-world scenarios and pick the best design for further physical testing and validation before delivering it to the customers.
Here is your task:
Pennsylvania’s fire department is using UAVs to monitor wildfires. The UAV should have great hover time (maintain target position with minimal deviation) and carry at least 10 kg (22.0 lbs) of monitoring equipment, including RGB cameras, IR cameras, and other sensors. The UAV must rise to a designated height at a fast vertical lift speed. Without sacrificing hover time, more carrying capacity is desirable for carrying more equipment for better coverage. The UAV must also fly steadily and sustain cross-wind and other potential environmental hazards for safety reasons. There is no requirement for maximum travel distance and travel speed.
Important Design Information: The best design refers to the design you deem most optimal considering all factors. Hover time is not air time. Hover time means the UAV is hovering at the target position in mid-air. The designs are not presented in any particular order. There is no correlation between the designs’ order and their performance. The designs are only presented in the order in which they are generated by the AI. This is an AI-generative system. The design may or may not work in real life. Please use your engineering knowledge and judgment. Consider all factors, including external ones, that may not have been considered and simulated by the AI.
Appendix B UAV Designs Used in the Study with Design Features
Table 5: Design Rendering Table - List 1 (2 options). * indicates the Pareto-TOPSIS optimal design.
List Design Rendering
List 1 - 2 options design_18393* [Uncaptioned image]
List 1 - 2 options design_16875 [Uncaptioned image]
Table 6: Design Rendering Table - List 2 (8 options). * indicates the Pareto-TOPSIS optimal design.
List Design Rendering List Design Rendering
List 2 - 8 options design_2986 [Uncaptioned image] List 2 - 8 options design_25944 [Uncaptioned image]
List 2 - 8 options design_9510 [Uncaptioned image] List 2 - 8 options design_27150 [Uncaptioned image]
List 2 - 8 options design_20155* [Uncaptioned image] List 2 - 8 options design_27450 [Uncaptioned image]
List 2 - 8 options design_25139 [Uncaptioned image] List 2 - 8 options design_27604 [Uncaptioned image]
Table 7: Design Rendering Table - List 3 (16 options). * indicates the Pareto-TOPSIS optimal design.
List Design Rendering List Design Rendering
List 3 - 16 options design_14962 [Uncaptioned image] List 3 - 16 options design_18802 [Uncaptioned image]
List 3 - 16 options design_15317 [Uncaptioned image] List 3 - 16 options design_18914 [Uncaptioned image]
List 3 - 16 options design_16763 [Uncaptioned image] List 3 - 16 options design_18952 [Uncaptioned image]
List 3 - 16 options design_18368 [Uncaptioned image] List 3 - 16 options design_19051 [Uncaptioned image]
Table 8: *
List Design Rendering List Design Rendering
List 3 - 16 options design_20320 [Uncaptioned image] List 3 - 16 options design_26003 [Uncaptioned image]
List 3 - 16 options design_20985* [Uncaptioned image] List 3 - 16 options design_26842 [Uncaptioned image]
List 3 - 16 options design_24246 [Uncaptioned image] List 3 - 16 options design_27369 [Uncaptioned image]
List 3 - 16 options design_25633 [Uncaptioned image] List 3 - 16 options design_27901 [Uncaptioned image]
Table 7 (Continued.) Design Rendering Table - List 3 (16 options) (Continued). * indicates the Pareto-TOPSIS optimal design.
Table 9: Design Features Table - Design Information Explicitly Available to the Participants. * indicates the Pareto-TOPSIS optimal design.
List
Design
Hover Time
Max Travel
Distance (meters)
Mass
Max Speed
($m/s$)
Battery
Voltage ($V$)
Total Cost
Max Lift
List 1 - 2 options design_18393* 428.67 7189.98 4.98 25 22.2 657.42 50.24
List 1 - 2 options design_16875 211.81 3477.01 4.82 22 11.1 844.66 27.09
List 2 - 8 options design_2986 246.01 2596.37 3.99 11 11.1 667.50 13.88
List 2 - 8 options design_9510 177.91 2627.52 3.63 18 11.1 482.14 10.72
List 2 - 8 options design_20155* 306.71 5866.26 6.24 21 22.2 995.12 45.00
List 2 - 8 options design_25139 181.70 2329.96 3.71 14 11.1 572.78 11.24
List 2 - 8 options design_25944 151.24 2004.73 3.15 14 11.1 638.42 9.61
List 2 - 8 options design_27150 203.33 2782.17 4.74 16 11.1 691.50 14.70
List 2 - 8 options design_27450 308.99 3001.63 3.75 10 11.1 622.88 10.79
List 2 - 8 options design_27604 230.78 2468.81 6.28 11 22.2 1453.79 18.42
List 3 - 16 options design_14962 207.87 3472.91 6.38 24 22.2 817.38 87.61
List 3 - 16 options design_15317 194.30 3096.53 4.70 30 22.2 806.02 154.35
List 3 - 16 options design_16763 209.68 3762.22 4.29 30 22.2 646.02 65.97
List 3 - 16 options design_18368 59.72 858.54 8.65 14 22.2 2218.34 81.42
List 3 - 16 options design_18802 145.98 2766.04 4.79 24 14.8 730.23 55.55
List 3 - 16 options design_18914 106.54 1592.01 8.85 15 11.1 1457.84 28.50
List 3 - 16 options design_18952 150.31 3144.53 6.93 22 22.2 880.04 41.56
List 3 - 16 options design_19051 191.56 3441.67 5.77 30 14.8 840.41 36.11
List 3 - 16 options design_20320 101.82 1998.72 7.29 36 22.2 1444.34 182.29
List 3 - 16 options design_20985* 237.54 6337.14 10.78 30 22.2 1583.72 241.37
List 3 - 16 options design_24246 189.96 3116.33 4.86 35 14.8 539.76 32.02
List 3 - 16 options design_25633 114.83 2133.10 5.15 40 22.2 1377.69 165.03
List 3 - 16 options design_26003 105.14 2282.49 7.89 24 11.1 1932.02 29.78
List 3 - 16 options design_26842 102.81 1760.31 3.87 29 11.1 850.34 52.65
List 3 - 16 options design_27369 195.16 4241.97 9.64 32 22.2 2299.70 95.79
List 3 - 16 options design_27901 109.52 2016.81 8.32 23 11.1 2450.76 32.90
Table 10: Design Features Table - Design Information NOT Explicitly Available to the Participants. * indicates the Pareto-TOPSIS optimal design.
Design
Max Thrust
Effective Lift
Propeller
Area ($m^{2}$)
Total Rod
Length (mmmm)
Rod Length to
Area Ratio ($m^{-1}$)
Number of
Rods
Number of
Connectors
Number of
Propellers
design_18393* 287.22 45.26 0.59 4172.23 7.13 12 8 4
design_16875 85.85 22.27 0.68 2988.41 4.37 6 0 6
design_2986 55.10 9.89 0.68 2562.89 3.75 6 0 6
design_9510 58.12 7.09 0.46 4363.65 9.58 12 8 4
design_20155* 218.41 38.76 0.43 2829.59 6.58 8 4 4
design_25139 57.50 7.53 0.49 1705.51 3.50 4 0 4
design_25944 43.18 6.46 0.31 1617.17 5.22 4 0 4
design_27150 79.16 9.96 0.68 4538.84 6.64 18 12 6
design_27450 71.03 7.04 0.68 2413.34 3.53 6 0 6
design_27604 200.79 12.14 0.52 1816.09 3.50 4 0 4
design_14962 388.49 81.23 0.78 5331.75 6.86 18 12 6
design_15317 538.61 149.65 0.88 2812.80 3.20 6 0 6
design_16763 319.27 61.68 0.46 2968.60 6.49 8 4 4
design_18368 171.79 72.77 0.60 2452.79 4.12 6 0 6
design_18802 128.99 50.76 0.68 3608.00 5.28 18 12 6
design_18914 223.89 19.65 1.69 7048.64 4.17 24 12 12
design_18952 137.83 34.63 0.51 4165.84 8.11 18 12 6
design_19051 231.15 30.34 0.81 3346.24 4.12 8 2 6
design_20320 643.47 175.00 0.88 6852.84 7.81 18 12 6
design_20985* 1000.72 230.59 0.88 6833.34 7.78 18 12 6
design_24246 250.13 27.16 0.59 5066.62 8.66 12 8 4
design_25633 271.33 159.88 0.19 2671.68 13.75 6 0 6
design_26003 171.97 21.89 0.78 5909.16 7.60 18 12 6
design_26842 105.69 48.78 0.52 1971.66 3.80 4 0 4
design_27369 138.71 86.15 0.51 4044.45 7.88 18 12 6
design_27901 249.50 24.58 1.17 3798.75 3.25 12 4 8
Table 11: Design Features Table - Identifying Features and Pick Percentage. * indicates the Pareto-TOPSIS optimal design.
Design
Identifying Feature
Non-
axisymetric?
Non-
planar?
Off-
plane?
Pick
% V
Pick
% D
Pick
% M
Pick
% Overall
design_18393* off-plane quadcopter NO NO YES 10.62% 96.88% 63.75% 57.08%
design_16875 hexacopter NO NO NO 89.38% 3.12% 36.25% 42.92%
design_2986 hexacopter NO NO NO 23.75% 8.12% 9.38% 13.75%
design_9510 off-plane quadcopter NO NO YES 1.25% 9.38% 13.12% 7.92%
design_20155* non-axisym non-planar off-plane quadcopter YES YES YES 1.25% 47.50% 30.00% 26.25%
design_25139 non-axisym quadcopter YES NO NO 6.25% 4.38% 5.00% 5.21%
design_25944 non-axisym quadcopter YES NO NO 3.75% 0.00% 1.25% 1.67%
design_27150 off-plane hexacopter NO NO YES 5.62% 5.00% 6.25% 5.62%
design_27450 hexacopter NO NO NO 16.88% 24.38% 31.88% 24.38%
design_27604 quadcopter NO NO NO 41.25% 1.25% 3.12% 15.21%
design_14962 off-plane hexacopter NO NO YES 0.62% 3.75% 3.12% 2.50%
design_15317 hexacopter NO NO NO 26.88% 8.75% 35.62% 23.75%
design_16763 non-axisym non-planar off-plane quadcopter YES YES YES 0.00% 23.75% 6.25% 10.00%
design_18368 hexacopter NO NO NO 16.25% 0.62% 0.62% 5.83%
design_18802 non-axisym non-planar off-plane hexacopter YES YES YES 1.25% 2.50% 1.25% 1.67%
design_18914 non-axisym non-planar off-plane YES YES YES 0.62% 0.62% 0.00% 0.42%
design_18952 non-axisym non-planar off-plane hexacopter YES YES YES 0.62% 0.00% 8.12% 2.92%
design_19051 non-axisym hexacopter YES NO NO 0.00% 11.25% 3.75% 5.00%
design_20320 off-plane hexacopter NO NO YES 0.00% 1.25% 0.00% 0.42%
design_20985* off-plane hexacopter NO NO YES 0.62% 41.25% 24.38% 22.08%
design_24246 off-plane quadcopter NO NO YES 0.00% 3.75% 3.12% 2.29%
design_25633 hexacopter NO NO NO 1.88% 0.00% 2.50% 1.46%
design_26003 non-axisym non-planar off-plane hexacopter YES YES YES 1.25% 0.00% 0.00% 0.42%
design_26842 quadcopter NO NO NO 33.12% 1.25% 6.25% 13.54%
design_27369 non-axisym non-planar off-plane hexacopter YES YES YES 6.25% 1.25% 4.38% 3.96%
design_27901 octocopter NO NO NO 10.62% 0.00% 0.62% 3.75%
Appendix C Survey Responses
C.1 Study 1
Experimental
Condition
Design A
(design 18393)
Design B
(design 16875)
Visual Rendering 4 (13.79%) 25 (86.21%)
Numerical Data 27 (93.10%) 2 (6.90%)
Visual Rendering
+ Numerical Data
17 (58.62%) 12 (41.38%)
Table 12: Study 1: The distributions of participants’ design choices for Design Problem 1 (2 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 2986)
Design B
(design 9510)
Design C
(design 20155)
Design D
(design 25139)
Design E
(design 25944)
Design F
(design 27150)
Design G
(design 27450)
Design H
(design 27604)
Visual 8
(27.59%)
1
(3.45%)
1
(3.45%)
0
(0.00%)
0
(0.00%)
0
(0.00%)
6
(20.69%)
13
(44.83%)
Data 3
(10.34%)
2
(6.90%)
13
(44.83%)
2
(6.90%)
0
(0.00%)
1
(3.45%)
7
(24.14%)
1
(3.45%)
Visual
+ Data
4
(13.79%)
6
(20.69%)
5
(17.24%)
1
(3.45%)
0
(0.00%)
0
(0.00%)
12
(41.38%)
1
(3.45%)
Table 13: Study 1: The distributions of participants’ design choices for Design Problem 2 (8 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 14962)
Design B
(design 15317)
Design C
(design 16763)
Design D
(design 18368)
Design E
(design 18802)
Design F
(design 18914)
Design G
(design 18952)
Design H
(design 19051)
Visual 0
(0.00%)
7 (24.14%) 0 (0.00%) 5 (17.24%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Data 1 (3.45%) 2 (6.90%) 12 (41.38%) 0 (0.00%) 1 (3.45%) 0 (0.00%) 0 (0.00%) 1 (3.45%)
Visual
+ Data
0 (0.00%) 13 (44.83%) 3 (10.34%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 2 (6.90%)
Exp.
Condition
Design I
(design 20320)
Design J
(design 20985)
Design K
(design 24246)
Design L
(design 25633)
Design M
(design 26003)
Design N
(design 26842)
Design O
(design 27369)
Design P
(design 27901)
Visual 0 (0.00%) 0 (0.00%) 0 (0.00%) 1 (3.45%) 0 (0.00%) 12 (41.38%) 0 (0.00%) 4 (13.79%)
Data 1 (3.45%) 10 (34.48%) 1 (3.45%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Visual
+ Data
0 (0.00%) 5 (17.24%) 0 (0.00%) 3 (10.34%) 0 (0.00%) 3 (10.34%) 0 (0.00%) 0 (0.00%)
Table 14: Study 1: The distributions of participants’ design choices for Design Problem 3 (16 options). The data is shown in the format of "number of participants (percentage)".
Experimental
Condition
Problem 1
(2 options)
Problem 2
(8 options)
Problem 3
(16 options)
Visual Rendering 4 (13.79%) 1 (3.45%) 0 (0.00%)
Numerical Data 27 (93.10%) 13 (44.83%) 10 (34.48%)
Visual Rendering
+ Numerical Data
17 (58.62%) 5 (17.24%) 5 (17.24%)
Table 15: Study 1: The number of participants who have chosen the optimal designs
C.2 Study 2
Experimental
Condition
Design A
(design 18393)
Design B
(design 16875)
Visual Rendering 13 (10.24%) 114 (89.76%)
Numerical Data 124 (97.64%) 3 (2.36%)
Visual Rendering
+ Numerical Data
84 (66.14%) 43 (33.86%)
Table 16: Study 2: The distributions of participants’ design choices for Design Problem 1 (2 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 2986)
Design B
(design 9510)
Design C
(design 20155)
Design D
(design 25139)
Design E
(design 25944)
Design F
(design 27150)
Design G
(design 27450)
Design H
(design 27604)
Visual 28
(22.05%)
1
(0.79%)
1
(0.79%)
10
(7.87%)
6
(4.72%)
8
(6.30%)
20
(15.75%)
53
(41.73%)
Data 10
(7.87%)
13
(10.24%)
60
(47.24%)
5
(3.94%)
0
(0.00%)
6
(4.72%)
32
(25.20%)
1
(0.79%)
Visual
+ Data
10
(7.87%)
15
(11.81%)
43
(33.86%)
7
(5.51%)
2
(1.57%)
8
(6.30%)
38
(29.92%)
4
(3.15%)
Table 17: Study 2: The distributions of participants’ design choices for Design Problem 2 (8 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 14962)
Design B
(design 15317)
Design C
(design 16763)
Design D
(design 18368)
Design E
(design 18802)
Design F
(design 18914)
Design G
(design 18952)
Design H
(design 19051)
Visual 1
(0.79%)
34
(26.77%)
0
(0.00%)
20
(15.75%)
1
(0.79%)
1
(0.79%)
1
(0.79%)
0
(0.00%)
Data 5
(3.94%)
11
(8.66%)
25
(19.68%)
1
(0.79%)
3
(2.36%)
1
(0.79%)
0
(0.00%)
16
(12.60%)
Visual
+ Data
5
(3.94%)
40
(31.50%)
7
(5.51%)
1
(0.79%)
2
(1.57%)
0
(0.00%)
13
(10.24%)
4
(3.15%)
Exp.
Condition
Design I
(design 20320)
Design J
(design 20985)
Design K
(design 24246)
Design L
(design 25633)
Design M
(design 26003)
Design N
(design 26842)
Design O
(design 27369)
Design P
(design 27901)
Visual 0
(0.00%)
1
(0.79%)
0
(0.00%)
2
(1.57%)
2
(1.57%)
41
(32.28%)
10
(7.87%)
13
(10.24%)
Data 1
(0.79%)
55
(43.31%)
5
(3.94%)
0
(0.00%)
0
(0.00%)
2
(1.57%)
2
(1.57%)
0
(0.00%)
Visual
+ Data
0
(0.00%)
34
(26.77%)
5
(3.94%)
1
(0.79%)
0
(0.00%)
7
(5.51%)
7
(5.51%)
1
(0.79%)
Table 18: Study 2: The distributions of participants’ design choices for Design Problem 3 (16 options). The data is shown in the format of "number of participants (percentage)".
Experimental
Condition
Problem 1
(2 options)
Problem 2
(8 options)
Problem 3
(16 options)
Visual Rendering 13 (10.24%) 1 (0.79%) 1 (0.79%)
Numerical Data 124 (97.64%) 60 (47.24%) 55 (43.31%)
Visual Rendering
+ Numerical Data
84 (66.14%) 43 (33.86%) 34 (26.77%)
Table 19: Study 2: The number of participants who have chosen the optimal designs
C.3 Study 3
Experimental
Condition
Design A
(design 18393)
Design B
(design 16875)
Visual Rendering 0 (0.00%) 4 (100.00%)
Numerical Data 4 (100.00%) 0 (0.00%)
Visual Rendering
+ Numerical Data
1 (25.00%) 3 (75.00%)
Table 20: Study 3: The distributions of participants’ design choices for Design Problem 1 (2 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 2986)
Design B
(design 9510)
Design C
(design 20155)
Design D
(design 25139)
Design E
(design 25944)
Design F
(design 27150)
Design G
(design 27450)
Design H
(design 27604)
Visual 2 (50.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 1 (25.00%) 1 (25.00%) 0 (0.00%)
Data 0 (0.00%) 0 (0.00%) 3 (75.00%) 0 (0.00%) 0 (0.00%) 1 (25.00%) 0 (0.00%) 0 (0.00%)
Visual
+ Data
1 (25.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 2 (50.00%) 1 (25.00%) 0 (0.00%)
Table 21: Study 3: The distributions of participants’ design choices for Design Problem 2 (8 options). The data is shown in the format of "number of participants (percentage)".
Exp.
Condition
Design A
(design 14962)
Design B
(design 15317)
Design C
(design 16763)
Design D
(design 18368)
Design E
(design 18802)
Design F
(design 18914)
Design G
(design 18952)
Design H
(design 19051)
Visual 0 (0.00%) 2 (50.00%) 0 (0.00%) 1 (25.00%) 1 (25.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Data 0 (0.00%) 1 (25.00%) 1 (25.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 1 (25.00%)
Visual
+ Data
0 (0.00%) 4 (100.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Exp.
Condition
Design I
(design 20320)
Design J
(design 20985)
Design K
(design 24246)
Design L
(design 25633)
Design M
(design 26003)
Design N
(design 26842)
Design O
(design 27369)
Design P
(design 27901)
Visual 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Data 0 (0.00%) 1 (25.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Visual
+ Data
0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Table 22: Study 3: The distributions of participants’ design choices for Design Problem 3 (16 options). The data is shown in the format of "number of participants (percentage)".
Experimental
Condition
Problem 1
(2 options)
Problem 2
(8 options)
Problem 3
(16 options)
Visual Rendering 0 (0.00%) 0 (0.00%) 0 (0.00%)
Numerical Data 4 (100.00%) 3 (75.00%) 1 (25.00%)
Visual Rendering
+ Numerical Data
1 (25.00%) 0 (0.00%) 0 (0.00%)
Table 23: Study 3: The number of participants who have chosen the optimal designs